# Hyperparameter Optimization in Machine Learning

### v2

@@ Abstract
Hyperparameters are configuration variables controlling the behavior of machine learning algorithms. They are ubiquitous in machine learning and artificial intelligence and the choice of their values determines the effectiveness of systems based on these technologies. Manual hyperparameter search is often unsatisfactory and becomes infeasible when the number of hyperparameters is large. Automating the search is an important step towards advancing, streamlining, and systematizing machine learning, freeing researchers and practitioners alike from the burden of finding a good set of hyperparameters by trial and error. In this survey, we present a unified treatment of hyperparameter optimization, providing the reader with examples, insights into the state-of-the-art, and numerous links to further reading. We cover the main families of techniques to automate hyperparameter search, often referred to as hyperparameter optimization or tuning, including random and quasi-random search, bandit-, model-, population-, and gradient-based approaches. We further discuss extensions, including online, constrained, and multi-objective formulations, touch upon connections with other fields such as meta-learning and neural architecture search, and conclude with open questions and future research directions.
@@
Over the last decade, machine learning has established itself as a key driver of innovation in science and industry. The digitalization of human activities and interactions has led to the generation of data at an unprecedented scale. These data could be stored ever more cheaply thanks to advances in micro-electronics, as well as the advent of cloud computing. Progress in hardware and low-power chip design also led to an exponential increase of computing capabilities of respectively cloud infrastructure, notably graphical processing units (GPUs), and devices on the edge, such as mobile phones. These two trends, cheaper and better hardware for storage and compute, combined with strong recent advances in algorithms and architectures for deep learning, lay the foundation for the successful application of statistical machine learning in practice and its adoption in artificial intelligence. It enabled learning algorithms to sift through vast amounts of data to discover and extract patterns that can then be used to make data-driven decisions without or with limited human intervention.
@@
One notable illustration of the recent advances made in machine learning is AlphaGo [silver:2017:mastering_go], a computer program developed by machine learning experts at the London-based company DeepMind. AlphaGo won against one of the world champions of the game of Go, Lee Sedol, in 2016. This event generated considerable press coverage (and was even made into a movie) as it was believed at the time that no computer program could win a game of Go against a human before several decades to come. What is less known is that the success of AlphaGo critically depended on how a set of configuration parameters, called hyperparameters, were adjusted automatically by another computer program [chen:2018:bayesian_optimization_alphago]. This computer program relied on Bayesian optimization, a technique that sequentially predicts and then evaluates the performance of algorithms such as AlphaGo when it uses specific hyperparameters. Until a certain cost or time budget is exhausted, Bayesian optimization repeatedly trades off exploration and exploitation to determine what is the next most interesting set of hyperparameters to evaluate based on the system performances observed so far.
In practice, the generalization capabilities of any machine learning algorithm also depend on hyperparameters. Whereas conventional (shallow) algorithms such as support vector machines [shawe-taylor:2004:kernel_methods_pattern_analysis] need hand-engineered features to transform the raw data in a suitable format, deep neural networks learn in an “end-to-end” fashion by ingesting the raw data directly [lecun2015nature]. However, in both cases, the composition of the model (e.g., the choice of feature extractors or the neural network architecture) and the empirical loss minimizer depend on configuration parameters that are not adjusted by the algorithm itself. Ultimately, their choice contributes to determine how effectively the algorithm can learn a specific natural language understanding, computer vision, or speech recognition task.
@@ 1.1 A simple illustrative example
To illustrate the practical importance of carefully selecting the hyperparameters, we consider the problem of sentiment analysis. yogotama2015-emnlp15 studied the effect of hyperparameters in this context. More specifically, the authors framed the problem as a standard binary classification problem as is commonly done in the literature, where the task of the classifier is to predict whether the text expresses a negative or a positive sentiment. They compared simple logistic regression trained with stochastic gradient descent to convolutional neural networks, which achieved state-of-the-art results at the time of the publication. Table 1.1(a) shows the hyperparameters that were searched over. These included the choice of text features used (e.g., removal of stop words or not), the type of regularization (e.g., $\ell_{1}$ or $\ell_{2}$) and optimization algorithm parameters (e.g., convergence tolerance). The experimental results they reported on the Amazon electronics dataset are reproduced in Table 1.1(b). Interestingly, they showed that tuned logistic regression using a bag of words representation of the text performed on par with a convolutional neural network and only slightly worse than a sequential convolutional neural network. Hence, by carefully optimizing hyperparameters, the authors could achieve a result close to the state of the art with a much simpler model. We will come back to this example in more details in the next chapter.
@@
Figure 1.1: Results on sentiment analysis reported in [yogotama2015-emnlp15]. Hyperparameters that were optimized and associated domains are listed on the left table. The feature extraction depends on the range of $n$-grams considered through $n_{min}$ and $n_{max}$. Stop words could be removed or not and words weighted according different weighting schemes. Two choices of regularization were considered and the amount of regularization was optimized. The convergence tolerance was also optimized. The binary classifiers compared are either SVM (support vector machine), LR (logistic regression), NN (multi-layer perceptron) or CNN (convolutional neural network). The authors found that when optimizing all these hyperparameters, simple logistic regression trained by stochastic gradient descent was performing as well as convolutional neural networks, which were the state of the art at the time of publication.
Hyperparameters play a critical role in machine learning as they do not only determine the generalization capabilities of trained models, but they may actually determine what constitutes the state-of-the-art. Indeed, results reported in an empirical study might look very differently for another choice of hyperparameters and so would be the conclusions drawn. Unfortunately, it is not uncommon for published results to omit specific hyperparameters used in the experiments, which may hinder reproducibility. Moreover, when experiments support the claim that a proposed method surpasses previous ones, the instability of results with respect to hyperparameters may cast doubts on these claims and raise questions about the current state-of-the-art [haibe-kains:2020:transparency].
The subject of this monograph is hyperparameter optimization (HPO). The availability of algorithmic tools to automate this task and their adoption will help machine learning researchers dealing with hyperparameters in a principled way. Importantly, these tools will also significantly help towards ensuring that new published advances are reproducible and thus unequivocally contribute to the field. Along with the experimental set-up, we argue that researchers would need to publish the HPO algorithm they used, the amount of resources it was allowed to consume [dodge:2019:show_your_work], and the hyperparameter search space that was considered [gundersen:2018:state_art]. Our objective is to provide a comprehensive overview of established techniques and recent advances in HPO. By doing so we hope we will help establish the practice of systematically using automatic HPO algorithms in research and industry, besides offering an introductory guide to understanding the fundamentals of this field.
@@ 1.2 What is hyperparameter optimization?
Let $A(\lambda)$ denote a learning algorithm configured by the hyperparameter vector $\lambda$ and consider a function $\mathscr{M}$ that evaluates its performance on a learning task. We call the composition $f(\lambda)=\mathscr{M}\circ A(\lambda)$ the response function. HPO consists in optimizing $f$ with respect to $\lambda$.
What distinguishes the HPO problem from conventional optimization problems, such as those arising in convex or combinatorial optimization, is its nested nature: to evaluate the response function $f$ at a specific point $\lambda$, it is required to run the learning algorithm $A$, which in turn has to solve another optimization problem to fit a model to the available data. In practice, running $A$ can be time-consuming. For example, training ResNet-50 from scratch on CIFAR-10 on an Amazon Elastic Cloud Compute (EC2) g4dn.xlarge instance takes more than 2 hours. [^1] A ResNet is a deep residual neural network commonly used in computer vision. ResNet-50 has 50 convolutional layers with skip connections (one for every stack of three layers). We refer the interested reader to the original paper by he2016deep.
@@ 1.3 Why is hyperparameter optimization difficult?
In general, computing $f$ might not only be time-consuming, but also expensive in terms of amount of resources used or cost. Its analytical form is typically either unknown or impractical to handle, it might have multiple local optima, it might exhibit wide flat regions, or it might be even non-smooth or non-continuous as a function of $\lambda$. As we will see in the following, for some classes of algorithms, we only require that $f$ is observable. Returning to the example of AlphaGo, this means that it does not matter whether the mathematical expression linking the hyperparameters of AlphaGo (i.e., how the algorithm is configured) and the probability of winning a game is known; what matters is that, for our specific choice of $\lambda$, we can observe if AlphaGo won or lost the game. Therefore, in general, one cannot apply conventional gradient-based optimization algorithms to solve HPO problems (but we will see that with additional assumption on $f$ this is, in fact, possible).
In addition, evaluations of $f$ can be stochastic, and the search space for $\lambda$ can itself be quite complex. Some hyperparameters could be continuous (e.g., learning rate), but others could be integral (e.g., number of layers in a deep neural network) or categorical to encode choices. Even more structure can emerge in the search space: conditional hyperparameters are variables that are only relevant depending on the value taken by others. For example, the number of units of the $j$-th layer of a neural network is only defined if the network has at least $j$ layers. If the domain to search over would be a small discrete set, one could carry out an exhaustive search by evaluating all the possible hyperparameter configurations in this set. Unfortunately, this is rarely the case in reality as the number of hyperparameter configurations to consider can be large or even uncountable.
All these challenges mean that, in practice, the focus of HPO is more often about finding “reasonably good” configurations as quickly as possible — which could still mean days of computation — rather than finding (global) optima. Hence, HPO often consists in determining a small set of good hyperparameters on which to evaluate $f$.
@@ 1.4 Historical remarks
Modern machine learning often considers a hierarchy of model families, indexed by model parameters at the lower level and hyperparameters at the higher level. Traditional model selection, such as Akaike’s Information Criterion (AIC) and the Bayesian Information Criterion (BIC), focuses on choosing the “model size” (i.e., a single integer-valued hyperparameter that determines the number of model parameters) that results in a good generalization [akaike:1974:new_look_statistical, schwarz1978estimating]. Interestingly, the origins of some recent HPO methods can be traced back to at least two decades before the introduction of AIC and BIC: the HPO problem in machine learning can be seen as a special case of optimal experimental design, an area that has a long tradition in statistics since the introduction of the response surface methodology [box:1951:experimental_attainment_optimum]; see myers16:_respon for a recent account.
Many ideas useful for tuning hyperparameters originated from a different class of problems: the configuration of heuristics and meta-heuristics in problem solving in Computer Science. The goal in these cases is not generalization but finding a solution with a small amount of computational resources. For example the LEX system [mitchell:1983:learning_experimentation_acquiringa] already embodied some fundamental ideas behind model-based approaches to HPO. LEX ran the to-be-configured algorithm to produce examples from which a generalizer learned to produce better heuristics. Model-based approaches have been subsequently applied to global optimization and propositional satisfiability [hutter:2009:experimental_investigation_modelbased], and to supervised learning [thornton2013auto]. These model-based approaches are based on or closely related to Bayesian optimization [mockus:1975:bayesian_methods_seeking, sacks:1989:design_analysis_computer, jones1998efficient]; see also the review by shahriari:2016:taking_human_out. Bayesian optimization is a technique at the core of current model-based HPO approaches in machine learning (see Section 4.2). Yet another class of approaches is based on the idea of racing, where candidate configurations are run in parallel and, as soon as sufficient statistical evidence is collected against a candidate, the run is stopped and the candidate replaced by a new one. This idea was initially formulated for model selection in machine learning [maron:1994:hoeffding_races_accelerating], and then applied to configuration of meta-heuristics [birattari:2002:racing_algorithm_configuring]. Finally, population-based algorithms also originated in the area of problem solving with seminal works of rechenberg65 and fogel:1965:intelligent_decisionmaking_simulation, boosted by the introduction of genetic inspired operators by holland75. While their application to optimization in machine learning has a long tradition, the use of population-based algorithms in HPO is more recent and preceded by the idea of gradient-based approaches [bengio_gradient-based_2000].
@@ 1.5 Outline of the monograph
In this chapter, we gave a broad overview of the HPO problem, motivating its importance and its place in machine learning. In the next chapter, we will formalize many concept introduced here and further present several examples of hyperparameters and HPO problems. In Chapter 3, we will introduce elementary algorithms that search over a predefined or random grid of hyperparameters. While conceptually simple, random search strategies based on uniform sampling are attractive as they are embarrassingly parallel and exhibit an anytime behavior. However, they require the evaluation of a relatively large number hyperparameters to find a good solution. In Chapter 4, we will study model-based strategies, such as Bayesian optimization. These approaches learn a surrogate model of the response function $f(\lambda)$ based on the evaluations observed so far. Model-based approaches are sample-efficient, which means that they typically require a much smaller number of hyperparameters (compared to grid or random search) to be evaluated to find a satisfactory solution. However, they are inherently sequential, which can negatively impact the run time of the optimization algorithm. In Chapter 5, we will discuss how randomized approaches can be made resource-efficient by stopping unpromising evaluations early. We will also show that they can be married to model-based approaches in order to sample the hyperparameter search space in a non-uniform fashion. Chapter 6 covers population-based methods, stochastic algorithms for global optimization that draw inspirations from natural phenomena. When applied to HPO, some techniques in this family such as population-based training maintain a set of predictors whose hyperparameters are updated dynamically, unlike the aforementioned approaches where hyperparameters are fixed for the entire training duration. In Chapter 7, we will discuss how gradient-based approaches can be applied to the optimization of continuous hyperparameters through hypergradients, which is more time-efficient than search-based algorithms. Finally, we close the monograph by discussing advanced use cases. In particular, we study how to adapt HPO to constrained and multi-objective settings. Indeed, HPO algorithms provide a powerful set of tools to meet competing requirements in applications (e.g., maximizing the predictive performance while mitigating the prediction bias) and satisfy practical constraints (e.g., maximal prediction latency). Furthermore, we also discuss connections with with neural architecture search and meta-learning, and touch upon transfer learning of hyperparameters for foundation models. The two closing chapters are dedicated to an overview of HPO software frameworks (Chapter 9) and emerging topics and open questions in the field (Chapter 10).

### v3

@@ Abstract
Hyperparameters are configuration variables controlling the behavior of machine learning algorithms. They are ubiquitous in machine learning and artificial intelligence and the choice of their values determines the effectiveness of systems based on these technologies. Manual hyperparameter search is often time-consuming and becomes infeasible when the number of hyperparameters is large. Automating the search is an important step towards advancing, streamlining, and systematizing machine learning, freeing researchers and practitioners alike from the burden of finding a good set of hyperparameters by trial and error. In this survey, we present a unified treatment of hyperparameter optimization, providing the reader with examples, insights into the state-of-the-art, and numerous links to further reading. We cover the main families of techniques to automate hyperparameter search, often referred to as hyperparameter optimization or tuning, including random and quasi-random search, bandit-, model-, population-, and gradient-based approaches. We further discuss extensions, including online, constrained, and multi-objective formulations, touch upon connections with other fields, such as meta-learning and neural architecture search, and conclude with open questions and future research directions.
@@
In this chapter, we informally introduce the main subject of this monograph: algorithms for optimizing hyperparameters. Hyperparameters are configuration variables that are ubiquitous in machine learning methods and can strongly influence generalization and overall performance. The wide array of existing algorithms for optimizing hyperparameters is the product of several years of research. Their adoption is already widespread in academia and in the industry. Our goal is to provide a comprehensive introduction to the fundamentals of these methodologies, with the aim of encouraging their broader adoption – making them standard practice —- and providing a solid foundation for further research in the field. Such efforts help strengthen and advance the current state-of-the-art, foster transparency and improve the reproducibility of scientific and engineering results. This is a timely necessity as the application of AI has rapidly expanded during the last decade and has solidified its position as a primary driver of innovation in many areas of science and industry.
@@
The digitalization of human activities and interactions has led to the generation of data at an unprecedented scale. On the one hand, data can now be stored more affordably, thanks to advancements in microelectronics and the emergence of cloud computing. On the other hand, progress in hardware and low-power chip design have led to an exponential increase in computing capabilities for both cloud infrastructure, notably graphical processing units (GPUs), and on edge devices, like mobile phones. The convergence of these two key trends has been instrumental in the pervasive adoption of machine learning. Recent breakthroughs in algorithms and architectures for deep learning have further reinforced this foundation. Learning algorithms can now sift through vast amounts of data to discover and extract patterns that can then be used to guide decisions with little or even no human intervention. But this is largely due to a dramatic increase in model complexity, both in terms of size and structure. Modern architectures often comprise several intertwined modules, each introducing its own set of hyperparameters. Without systematic tools, configuring these hyperparameters jointly can be complex and overwhelming.
@@
One notable and often cited illustration of the recent advances made in machine learning is AlphaGo [silver:2017:mastering_go], a computer program developed at the London-based company DeepMind. AlphaGo won against one of the world champions of the game of Go, Lee Sedol, in 2016. This event generated considerable press coverage (and was even made into a movie) as it was believed at the time that no computer program could win a game of Go against a human before several decades to come. What is less known is that the success of AlphaGo critically depended on how a set of hyperparameters were adjusted automatically by another computer program [chen:2018:bayesian_optimization_alphago]. This computer program relied on Bayesian optimization (see Section 4.2), a general black-box technique that in this case sequentially predicts and then evaluates the performance of learning algorithms, such as those behind AlphaGo, when their hyperparameters are set to specific values.
@@ 1.1 A Simple Illustrative Example
To illustrate the practical importance of carefully selecting the hyperparameters, we consider the problem of sentiment analysis. yogotama2015-emnlp15 studied the effect of hyperparameters in this context. More specifically, the authors framed the problem as a standard binary classification problem as was commonly done in the literature, where the task of the classifier is to predict whether the text expresses a negative or a positive sentiment. They compared simple logistic regression, trained by stochastic gradient descent, to convolutional neural networks, which achieved state-of-the-art results at the time of the publication. Figure 1.1(a) shows the hyperparameters that were searched over. These included the choice of text features used (e.g., removal of stop words or not), the type of regularization (e.g., $L_{1}$ or $L_{2}$) and optimization algorithm parameters (e.g., convergence tolerance). The experimental results they reported on the Amazon electronics dataset are reproduced in Figure 1.1(b). Interestingly, they showed that an optimized logistic regression with a bag of words representation of the text performed similarly to a convolutional neural network and only slightly worse than a sequential convolutional neural network. Hence, by carefully optimizing the hyperparameters, the authors could achieve a result close to the state-of-the-art, yet using a much simpler model. In Section 2.1.1, we will come back to this example, explaining in detail the role of individual hyperparameters.
@@
Figure 1.1: Results on sentiment analysis reported in [yogotama2015-emnlp15]. Hyperparameters and their domains (both continuous and discrete) are listed on the left. Here $n_{min},n_{max}]$ denotes the $n$-gram range. SVM: support vector machine; LR: logistic regression; NN: multi-layer perceptron; CNN: convolutional neural network.
As this simple example reveals, claims about state-of-the-art performance improvements need to be interpreted carefully: empirical results can vary greatly depending on the chosen hyperparameters, with a possibly significant impact on the conclusions drawn. Unfortunately, it is not uncommon for published results to omit values for specific hyperparameters used in the experiments, or to omit details explaining how these values were chosen. In these cases, reproducibility may be hindered [haibe-kains:2020:transparency].
@@ 1.2 What Is Hyperparameter Optimization?
A typical supervised learning algorithm receives training data and outputs a predictor (e.g., a mapping from textual reviews to binary sentiments, as in our previous example). The quality of the predictor can then be measured on new data (the validation set) using an evaluation metric, such as the error rate. Since the predictor depends on the chosen hyperparameters, the validation error also depends on those hyperparameters. The mapping from hyperparameter values to the performance measured on the validation set is called the response function.
In a nutshell, hyperparameter optimization (HPO) consists in finding the hyperparameters that optimize the response function. What distinguishes the HPO problem from conventional optimization problems, such as those arising in convex or combinatorial optimization, is its nested nature: evaluating the response function at a specific point requires running the learning algorithm, which in turn often involves solving another optimization problem to fit a model to the available (training) data.
@@ 1.3 Why Is Hyperparameter Optimization Difficult?
First, the nested nature of hyperparameter optimization problems means that, in most cases, the response function is not available in closed form, but is defined only implicitly. The function may be stochastic (due to the inherent randomness of the training algorithm), have multiple local optima, exhibit wide flat regions, or be non-smooth, non-continuous, or non-differentiable. These characteristics rule out the application of standard gradient-based optimization algorithms. Moreover, evaluating the response function at a given hyperparameter assignment can be very expensive: for instance, training a modern deep architecture may require significant compute, ranging from several hours to several days on a GPU cluster. As a result, much of the field has focused on developing algorithms that make minimal assumptions, such as only requiring observability for instance. This perspective becomes clearer when we return to the example of AlphaGo: it does not matter whether the mathematical expression linking the hyperparameters of AlphaGo (i.e., how the algorithm is configured) to the probability of winning a game is explicitly known; what matters is that, for a given set of hyperparameters, we can observe whether AlphaGo won or lost the game.
Second, the hyperparameter search space is often complex and heterogeneous. Some hyperparameters could be continuous (e.g., the learning rate), but others could be integer-valued (e.g., the number of layers in a deep neural network) or categorical to encode choices. Additional structure can also arise in the search space. Notably, conditional hyperparameters are variables that are only relevant depending on the value taken by others. For example, the number of units in the $j$-th layer of a neural network is only relevant if the network has at least $j$ layers. If the domain to be searched is a small discrete set, an exhaustive search can be performed by evaluating all possible hyperparameter configurations within this set. Unfortunately, this is rarely the case. In most scenarios, the number of configurations can be infinite or even uncountable.
These challenges mean that, in practice, the focus of HPO is often to find “reasonably good” configurations as quickly as possible — which could still mean days of computation — rather than to find the global optimum of the response function. Hence, HPO consists in determining a small set of promising hyperparameters on which to evaluate the response function.
@@ 1.4 Historical Remarks
Although HPO has significantly developed in recent years, especially after the rise of deep learning, its foundations are much older. Here we briefly summarize some important landmarks. Model selection in statistics attracted significant attention in the 1970s and it is perhaps the earliest precursor to HPO in machine learning. Indeed, HPO itself is sometimes referred to as “model selection” – although the expression originally only denoted the simpler problem of finding the appropriate dimension of a model. Unlike modern approaches to HPO, which are based on a validation set, early model selection methods, such as Akaike’s information criterion [akaike:1974:new_look_statistical] and the Bayesian information criterion [schwarz1978estimating], were based on the idea of penalizing model complexity by adding a term to the likelihood function. In the context of small, analytically treatable model, penalties make sense as they avoid always selecting the largest (best fitting) model [schwarz1978estimating]. A closely related idea, rooted in information theory, is the minimum description length principle [rissanen1978modeling], where model complexity is measured by the number of bits required to encode it.
Interestingly, some algorithmic ideas behind modern HPO methods are even older and can be traced back to well before the 1970s. The HPO problem in machine learning is related to optimal experimental design, an area that has a long tradition in statistics since the introduction of the response surface methodology [box:1951:experimental_attainment_optimum]; see [myers16:_respon] for a recent account. One of the core ideas in that context is the approximation, by means of a surrogate model, of an otherwise unknown true response function that needs to be optimized. Model-based algorithms for HPO share the very same idea and the HPO objective function draws its name from these methodologies.
Other ideas behind modern HPO originate from a community at the intersection between AI and operations research, interested in a different class of problems: the configuration of heuristics and meta-heuristics in problem solving. The goal in these cases is not generalization as in machine learning, but finding solutions quickly, for example to certain instances of (weighted) Boolean satisfiability. The LEX system [mitchell:1983:learning_experimentation_acquiringa] already embodied some ideas behind model-based approaches to HPO, which we cover in Chapter 4. LEX ran the to-be-configured algorithm to produce examples from which a generalizer is trained to produce better heuristics. Later, sequential model-based configuration (SMAC) was proposed to solve global optimization and propositional satisfiability problems [hutter:2009:experimental_investigation_modelbased, hutter_sequential_2011]. It was eventually applied to supervised learning [snoek_practical_2012, thornton2013auto]. SMAC is related to Bayesian optimization [mockus:1975:bayesian_methods_seeking, sacks:1989:design_analysis_computer, jones1998efficient] as it is one of the techniques at the core of current model-based HPO approaches.
Yet another class of early approaches is based on the idea of racing, where candidate configurations are run in parallel and, as soon as sufficient statistical evidence is collected against a candidate, the run is stopped and the candidate replaced by a new one. This idea was initially formulated for model selection in machine learning [maron:1994:hoeffding_races_accelerating], and then applied to configuration of meta-heuristics [birattari:2002:racing_algorithm_configuring]. Modern multi-fidelity strategies for HPO also discard unpromising candidates but in a more general setting (see Chapter 5).
Finally, population-based algorithms also originated in the area of problem solving with seminal works of rechenberg65 and fogel:1965:intelligent_decisionmaking_simulation, boosted by the introduction of genetic inspired operators by holland75. While their application to optimization in machine learning has a long tradition, the use of population-based algorithms in HPO is more recent and preceded by the idea of gradient-based approaches [bengio_gradient-based_2000].
@@ 1.5 Outline of the Monograph
In this chapter, we introduced the HPO problem, motivating its importance and its place in machine learning. In the next chapter, we will formalize many concept introduced here and further present several examples of hyperparameters and HPO problems. In Chapter 3, we will introduce elementary algorithms that search over a predefined or random grid of hyperparameters. While conceptually simple, random search strategies based on uniform sampling are attractive as they are embarrassingly parallel and exhibit an anytime behavior, meaning that stopping the search at any time yields a valid experiment. However, they require the evaluation of a relatively large number hyperparameters to find a good solution. In Chapter 4, we will study model-based strategies, such as Bayesian optimization. These approaches learn a surrogate model of the response function based on the evaluations observed so far. Model-based approaches are sample-efficient, which means that they typically require evaluating a much smaller number of hyperparameters (compared to grid or random search) to find a satisfactory solution. However, they are inherently sequential, which can negatively impact the run time of the optimization algorithm. In Chapter 5, we will discuss how randomized approaches can be made resource-efficient by stopping unpromising evaluations early. We will also show that they can be married to model-based approaches in order to sample the hyperparameter search space in a non-uniform fashion. Chapter 6 covers population-based methods, stochastic algorithms for global optimization that draw inspirations from natural phenomena. When applied to HPO, some techniques in this family, such as population-based training, maintain a set of predictors whose hyperparameters are updated dynamically, unlike the aforementioned approaches where hyperparameters are fixed for the entire training duration. In Chapter 7, we will discuss how gradient-based approaches can be applied to the optimization of continuous hyperparameters through hypergradients, which is more time-efficient than search-based algorithms. Finally, we close the monograph by discussing advanced use cases. In particular, we study how to adapt HPO to constrained and multi-objective settings. Indeed, HPO algorithms provide a powerful set of tools to meet competing requirements in applications (e.g., maximizing the predictive performance while mitigating the prediction bias) and satisfy practical constraints (e.g., maximal prediction latency). We also discuss connections with with neural architecture search and meta-learning, and touch upon transfer learning of hyperparameters for foundation models. The two closing chapters are dedicated to an overview of HPO software frameworks (Chapter 9) and emerging topics and open questions in the field (Chapter 10).